OpenStudy (anonymous):

What are the complex roots of X^3 + X^2 + X + 1 Express them both in standard form and in exponential notation

OpenStudy (mertsj):

\[x ^{3}+x ^{2}+x+1=(x+1)(x ^{2}+1)\]

OpenStudy (mertsj):

So the real root is -1 and the complex roots are positive or negative i or positive or negative (-1)^1/2

OpenStudy (mertsj):

\[x ^{2}=-1, x=\pm(i), or,\pm \sqrt{-1}\]

OpenStudy (mertsj):

\[\pm \sqrt{-1}= \pm(-1)^{-\frac{1}{2}}\]

OpenStudy (anonymous):

thanks!